It is of interest to decide if an analytical separation of the metal ions can be effected by selective precipitation of carbonates from a solution that is 9.11×10-2 M in Cu2+ and 0.100 M in Ca2+.

CaCO3 Ksp = 3.80×10-9
CuCO3 Ksp = 2.50×10-10


To analyze this problem, answer the following questions.

(1) What carbonate concentration is needed to precipitate 99.9% of the metal that forms the least soluble carbonate? __M

(2) When 99.9% of the least soluble carbonate has precipitated, will all of the metal that forms the more soluble carbonate still remain in solution? Yes/No?

(3) What is the upper limit on the carbonate ion concentration if the more soluble compound is not to precipitate? __M

(4) If the [CO32-] is at this upper limit, what percentage of the metal that forms the least soluble carbonate remains in solution? __%

What is the equilibrium Mg2+ concentration when 1.89 L of a 0.162 M magnesium nitrate solution are mixed with 2.04 L of a 0.282 M ammonium carbonate solution?

[Mg2+] = __M
Answer #1

(1) To solve this we have to take the Kps from least soluble salt, as we can realize the least soluble is CuCO3 , because it has a lower Kps

knowing that Kps = [Cu+2][CO3-2] = 2.50 x10-10

and now we want to precipitate the 99.9% of the Cu, it mean that the 0.1% of Cu will stay in solution, so we calculate the Cu that will stay in solution

[Cu+2] = (0.1/100) * (9.11 x10-2) = 9.11 x10-5 M

then the carbonate needed would be given by [CO3-2] =Kps / [Cu+2]

[CO3-2] = 2.50 x10-10 / 9.11 x10-5 = 2.7442 x10-6

(2) to know if the another coumpound will stay in solution, we could call Q , the product of the concentration of the ions, so we make

Q = [Ca+2][CO3-2] = (0.1) * ( 2.7442 x10-6) = 2.7442 x10-7   

Q > Kps this means it will also precipitate

(3) to know solubility, the upper limit, we make this change [Cu+2] = [CO3-2] = S

so Kps expresion would be

Kps = S2 and this means S = (Kps)1/2 = (2.50 x10-10)1/2 = 1.5811 x10-5

(4) we will use the Kps formula, but now we will calculate  [Cu+2]

this is

[Cu+2] =Kps / [CO3-2]

using the new concentration of cabonate, we get

[Cu+2] = 2.50 x10-10 / 1.5811 x10-5 = 1.5811 x10-5

so we would get (1 - 1.5811 x10-5 /9.11 x10-2) * 100% = 99.98%

(5)

to know the concentration of [Mg+2] , we start calculating the new concentration of all compounds

[Mg(NO3)2] = 1.89 * 0.162 / (1.89 +2.04) = 7.7908 x10-2 M

[(NH4)2CO3] = 2.04 * 0.282 / (2.04 + 1.89) = 0.1464 M

we can notice tha there is only an atom of Mg in every molecule of [Mg(NO3)2] and also onle a molecule of CO3 in every molecule of [(NH4)2CO3], also we know that they will be those that will precipitate, so let's calculate Q as following, using the concentrations calculated and know the presence of the Mg and CO3.

Q = [Mg+2] [CO3-2] = (7.7908 x10-2 ) * ( 0.1464) = 0.01140

The Kps of this salt is 6.82 x10-6 , Q is greater than Kps, it means that in the solution will only stay the solubility that can have the solution, calculated by

S = (Kps)1/2 = (6.82 x10-6 )1/2 = 2.6115 x10-3

Finally we could say that the [MG+2] that will stay in solution will be 2.6115 x10-3 M